# Whirling injections, surjections, and other functions between finite sets

### v3

@@ 3. Whirling restricted growth words
A partition of $[n]$ is said to be noncrossing if whenever $i<j<k<\ell$, it is not the case that $i$ and $k$ belong to one block of $\pi$ with $j$ and $\ell$ belonging to another block. A noncrossing RG-word is an RG-word whose corresponding partition is noncrossing. Let $\operatorname{RG}_{\mathrm{nc}}(n,k)$ and $\operatorname{RG}_{\mathrm{nc}}(n)$ denote the sets of all noncrossing RG-words in $\operatorname{RG}(n,k)$ and $\operatorname{RG}(n)$ respectively.
@@
An RG-word $f$ is noncrossing if and only if in the one-line notation, we do not have $abab$ in order (not necessarily consecutively) for some $a\not=b$.
@@
$123244123442\stackrel{{\scriptstyle w_{1}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{2}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{3}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{4}}}{{\longmapsto}}123242\stackrel{{\scriptstyle w_{5}}}{{\longmapsto}}123242\stackrel{{\scriptstyle w_{6}}}{{\longmapsto}}$ 123244
@@
For $\mathcal{F}=\operatorname{RG}(6)$,
$123224123442\stackrel{{\scriptstyle w_{1}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{2}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{3}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{4}}}{{\longmapsto}}123242\stackrel{{\scriptstyle w_{5}}}{{\longmapsto}}123222\stackrel{{\scriptstyle w_{6}}}{{\longmapsto}}$ 123224
@@
We conjecture homomesy for whirling noncrossing RG-words. Again $w_{1}$ acts trivially for $\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(n)$ or $\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(n,k)$, so for simplicity we ignore $w_{1}$.
@@
Let $n\geq 2$. Fix $\mathcal{F}$ to be either $\operatorname{RG}_{\mathrm{nc}}(n)$ or $\operatorname{RG}_{\mathrm{nc}}(n,k)$ for some $1\leq k\leq n$. Let $\pi$ be a permutation of $\{2,3,\dots,n\}$. Under the action of $\mathbf{w}_{\pi}$ on $\mathcal{F}$, $I_{2\mapsto 1}-I_{n\mapsto 1}$ is 0-mesic.
The proof of Theorem 3.8 relied on the fact from Lemma 3.12 that for $\mathcal{F}=\operatorname{RG}(n,k)$ or $\mathcal{F}=\operatorname{RG}(n)$, #$\#\{f\in\mathcal{F}\;|\;f(i)=1\}=\#\{f\in\mathcal{F}\;|\;f(j)=1\}$ given any $2\leq i,j\leq n$. For $\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(n,k)$ or $\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(n)$, we only conjecture this when $i=2$, $j=n$. For example, over the set of noncrossing partitions on $[4]$, there are five partitions containing 1 and 2 in the same block, four partitions containing 1 and 3 in the same block, and five partitions containing 1 and 4 in the same block.
Under the whirling left-to-right order ($\pi$ is the identity) Conjecture 3.18 has been confirmed for all $n\leq 9$ and relevant $k$ values ($k\in[n]$). It has also been tested and confirmed for many random whirling orders (with $n\leq 9$).

### v4

@@ 3.3. Whirling on noncrossing RG-words
A partition of $[n]$ is said to be noncrossing if whenever $i<j<k<\ell$, it is not the case that $i$ and $k$ belong to one block of $\pi$ with $j$ and $\ell$ belonging to another block. It is well-known that the number of noncrossing partitions of $[n]$ is given by the Catalan number $C_{n}=\frac{1}{n+1}\binom{2n}{n}$ and that the number of noncrossing partitions of $[n]$ into $k$ blocks is given by the Narayana number $N(n,k)=\frac{1}{n}\binom{n}{k}\binom{n}{k-1}$. A noncrossing RG-word is an RG-word whose corresponding partition is noncrossing. Let $\operatorname{RG}_{\mathrm{nc}}(n,k)$ and $\operatorname{RG}_{\mathrm{nc}}(n)$ denote the sets of all noncrossing RG-words in $\operatorname{RG}(n,k)$ and $\operatorname{RG}(n)$ respectively.
@@
An RG-word $f$ is noncrossing if and only if in the one-line notation, we do not have abababab in order (not necessarily consecutively) for some $a\not=b$.
@@
$123442\stackrel{{\scriptstyle w_{1}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{2}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{3}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{4}}}{{\longmapsto}}123242\stackrel{{\scriptstyle w_{5}}}{{\longmapsto}}123242\stackrel{{\scriptstyle w_{6}}}{{\longmapsto}}123244$
@@
For $\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(6)$,
$123442\stackrel{{\scriptstyle w_{1}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{2}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{3}}}{{\longmapsto}}123442\stackrel{{\scriptstyle w_{4}}}{{\longmapsto}}123242\stackrel{{\scriptstyle w_{5}}}{{\longmapsto}}123222\stackrel{{\scriptstyle w_{6}}}{{\longmapsto}}123224$
@@
Theorem 3.18.
Let $\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(n,k)$ for some $1\leq k\leq n$. Under the action of $\mathbf{w}$ on $\mathcal{F}$, $I_{2\mapsto 1}-I_{n\mapsto 1}$ is 0-mesic.
We also conjecture that this homomesy holds for other whirling orders, and on $\mathcal{F}=\operatorname{RG}(n)$.
@@
Fix $\mathcal{F}$ to be either $\operatorname{RG}_{\mathrm{nc}}(n)$ or $\operatorname{RG}_{\mathrm{nc}}(n,k)$ for some $1\leq k\leq n$. Let $\pi$ be a permutation of $\{2,3,\dots,n\}$. Under the action of $\mathbf{w}_{\pi}$ on $\mathcal{F}$, $I_{2\mapsto 1}-I_{n\mapsto 1}$ is 0-mesic.
We utilize some lemmas in the proof of Theorem 3.18.
Lemma 3.20.
Let $f\in\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(n,k)$ for some $1\leq k\leq n$, and let $1<i<n$. If f(i)=f(i1)f(i)=f(i-1), then $(w_{i}(f))(i)=f(i+1)$.
Proof.
Let a=f(i1)=f(i)a=f(i-1)=f(i) and $b=f(i+1)$. Clearly we can change the value of $f(i)$ from $a$ to $b$, and it will remain a noncrossing RG-word. We will now explain why changing the value of $f(i)$ to the various other values $c$ that we encounter on the way from $a$ to $b$ when we apply $w_{i}$ will not yield have a noncrossing RG-word.
Case 1: $a=b$. Since $f\in\operatorname{RG}_{\mathrm{nc}}(n,k)$, there must be some $j$ for which $f(j)=c$. So changing the value of $f(i)$ to $c$ will not yield a noncrossing RG-word by Proposition 3.15.
Case 2: $a<c<b$. Since $f$ is an RG-word, there must exist j<h<i1j<h<i-1 such that $f(j)=a$ and $f(h)=c$. So changing the value of $f(i)$ to $c$ will not yield a noncrossing RG-word by Proposition 3.15.
Case 3: $b<a<c$. Since $f\in\operatorname{RG}_{\mathrm{nc}}(n,k)$, there must be some $j$ for which $f(j)=c$. If j<i1j<i-1, then since $f$ is an RG-word, there must exist $h<j$ for which $f(h)=a$. Then changing the value of $f(i)$ to $c$ will not yield a noncrossing RG-word by Proposition 3.15. On the other hand, suppose $j>i+1$. Since $f$ is an RG-word, there must exist h<i1h<i-1 for which $f(h)=b$, in which case changing the value of $f(i)$ to $c$ will not yield a noncrossing RG-word by Proposition 3.15.
Case 4: $c<b<a$. Since $f$ is an RG-word, we must have h<j<i1h<j<i-1 satisfying $f(h)=c$ and $f(j)=b$. So changing the value of $f(i)$ to $c$ will not yield a noncrossing RG-word by Proposition 3.15. ∎
Lemma 3.21.
Let $f\in\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(n,k)$ for some $1\leq k\leq n$. If f(n1)=f(n)f(n-1)=f(n), then $(w_{n}(f))(i)=1$.
Proof.
Let a=f(n1)a=f(n-1). Let $a<b\leq k$. Since $f\in\operatorname{RG}_{\mathrm{nc}}(n,k)$, there must exist i<j<n1i<j<n-1 satisfying $f(i)=a$ and $f(j)=b$. So we cannot change the value of $f(n)$ to $b$ and still have a noncrossing RG-word by Proposition 3.15. However, changing the value of $f(n)$ to 1 will still yield a noncrossing RG-word. So $w_{n}$ changes the value of $f(n)$ to 1. ∎
Lemma 3.22.
Let $\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(n,k)$ for some $1\leq k\leq n$. Let $f\in\mathcal{F}$ for which $f(2)=1$. Then $(\mathbf{w}(f))(n)=1$ and $(\mathbf{w}(f))(i)=f(i+1)$ for all $2\leq i<n$.
The proof of the above lemma follows easily by Lemmas 3.20 and 3.21. We are now ready to prove Theorem 3.18.
Proof of Theorem 3.18.
Note that among noncrossing partitions of $[n]$ into $k$ blocks:
  • there are N(n1,k)N(n-1,k) of them that have 1 and 2 in the same block, and
  • there are N(n1,k)N(n-1,k) of them that have 1 and $n$ in the same block.
The first is because we can take any noncrossing partition of $\{2,3,\ldots,n\}$ and place 1 in the same block as 2 and still have a noncrossing partition. Likewise, the second is because we can take any noncrossing partition of [n1][n-1] and place $n$ in the same block as 1 and still have a noncrossing partition. So $\mathcal{F}$ contains the same number of functions $f$ satisfying $f(2)=1$ as functions $f$ satisfying $f(n)=1$. By Lemma 3.22, $\mathbf{w}$ is a bijection between functions $f$ satisfying $f(2)=1$ and functions $f$ satisfying $f(n)=1$. So each $\mathbf{w}$-orbit contains equal numbers of these two types of functions, and thus $I_{2\mapsto 1}-I_{n\mapsto 1}$ is 0-mesic. ∎